### v1

# Towards Interpretable Deep Neural Networks for Tabular Data

Abstract
Tabular data is the foundation of many applications in fields such as finance and healthcare. Although DNNs tailored for tabular data achieve competitive predictive performance, they are blackboxes with little interpretability. We introduce XNNTab, a neural architecture that uses a sparse autoencoder (SAE) to learn a dictionary of monosemantic features within the latent space used for prediction. Using an automated method, we assign human-interpretable semantics to these features. This allows us to represent predictions as linear combinations of semantically meaningful components. Empirical evaluations demonstrate that XNNTab attains performance on par with or exceeding that of state-of-the-art, black-box neural models and classical machine learning approaches while being fully interpretable.
1 Introduction
Tabular data is the most common type of data in a wide range of industries, including advertising, finance and healthcare. While deep neural networks (DNNs) achieved major advances in computer vision, their performance on tabular data remained below that of Gradient Boosted Decision Trees (GBDTs) (Chen and Guestrin, 2016; Khan et al., 2022). To improve the performance of DNNs on tabular data, several deep learning methods have been developed with specialized architectures that take into account the unique traits of tabular data (mixture of categorical, ordinal and continuous features) in order to be on par (Huang et al., 2020) or even outperform GBDTs (Chen et al., 2024). While these methods addressed the limitations of DNNs in terms of performance, DNNs still remain blackboxes, whose decisions are hard to communicate to end users (Schwalbe and Finzel, 2023).
In this paper, we address the interpretability limitation of existing DNNs and present XNNTab, a deep neural network with a sparse autoencoder (SAE) component. The SAE learns a dictionary of monosemantic, sparse features that are used for outcome prediction. We take inspiration from the success of SAEs in learning interpretable features from LLM’s internal activations (Huben et al., 2023; Yun et al., 2021). We use an automated method to assign human-understandable semantics to the learned monosemantic dictionary features. Our method is illustrated in Figure 1 and described in Section 3. We show that XNNTab outperforms both interpretable and classical blackbox models and has similar performance to GBDTs. Additionally, XNNTab is fully interpretable: the final prediction is a simple linear combination of the interpretable dictionary features, which have easily accessible human-understandable semantics.
2 Related Work
Gradient boosted decision trees (GBDTs) (Chen and Guestrin, 2016; Khan et al., 2022) have shown state-of-the art performance on tabular data (Grinsztajn et al., 2022). To close this performance gap to GBDTs, neural architectures specifically designed for tabular data have been proposed, such as attention-based architectures (Gorishniy et al., 2021; Somepalli et al., 2022; Yan et al., 2023), retrieval-augmented models (Somepalli et al., 2022; Gorishniy et al., 2023). Additional techniques include tailored regularizations (Jeffares et al., 2023), deep ensembles (Gorishniy et al., 2024), attentive feature selection strategies (Arik and Pfister, 2021), categorical feature embedding (Huang et al., 2020) and reconstruction of binned feature indices (Lee et al., 2024).
Recent methods focus not only on performance, but also aim to include interpretability. ProtoGate (Jiang et al., 2024) is an ante-hoc interpretable architecture for tabular data, which learns prototypes of classes and predicts new instances based on their similarity to class prototypes. InterpreTabNet (Si et al., 2024), a variant of TabNet (Arik and Pfister, 2021) adds post-hoc explanations by learning sparse feature attribution masks and using LLMs to interpret the learned features from the masks. Similar to ProtoGate and different to InterpetTabNet, XNNTab is intrinsically interpretable, i.e., it uses interpretable features directly for prediction. While ProtoGate relies on class prototypes, XNNTab learns relevant feature combinations that describe parts of instances, similar to part-prototype models (Elhadri et al., 2025).
3 Method
XNNTab consists of a standard neural blackbox (e.g., an MLP) to learn nonlinear latent features, and a sparse autoencoder (SAE) (Huben et al., 2023) to decompose this latent polysemantic feature space into monosemantic features (cf. Figure 1, A). The monosemantic features are in a latent space with unknown (but unique) semantics. To assign meaning to each feature, we learn rules that describe the subset of the training samples that highly activate each feature (cf. Figure 1, B). Finally, we combine the last linear layers into a single linear layer by multiplying their weight matrices (cf. Figure 1, C).
Figure 1: Architecture overview. A. An MLP learns nonlinear features, which are decomposed into monosemantic dictionary features using an SAE. B. Dictionary features are assigned a human-interpretable meaning by learning rules for the subset of training instances that highly activate a specific feature. C. Predictions are linear combinations of monosemantic features by combining the linear model components (the decoder, $M^{T}$, and the linear layer of the MLP, $W$).
3.1 Architecture
The base blackbox model is an MLP and with multiple hidden layers and a softmax output. We denote the penultimate layer as $l$, and the representations of this layer as $h_{l}$. The MLP learns a function $\hat{y}=f(x)$. $f(x)$ can be decomposed into a function $g$ learning the hidden representations $h_{l}$ and a linear model $h$ predicting the target based on $h_{l}$, i.e., $\hat{y}=h(h_{l})=h(g(x))$. $g(\cdot)$ is characterized by the parameters $\theta_{g}$, and $h(\cdot)$ is characterized by the weight matrix $W$.
For the SAE, we use the architecture introduced by Huben et al. (2023). The SAE consists of an input layer, a hidden layer and an output layer. The size of the hidden layer is $d_{hid}=R\cdot d_{in}$ with $d_{in}$ being the dimension of the MLP’s layer $l$, and $R$ a hyperparameter controlling the size of the feature dictionary. The input to the SAE are the hidden representations of the blackbox MLP $h_{l}$. The weights of the encoder and decoder are tied, and given by the matrix $M$ ($M^{T}$ for the decoder)  [^1] Additionally, the encoder has a bias vector $b$.. The output of the SAE is given by $\hat{h}_{l}=M^{T}\Bigl{(}RELU(Mh_{l}+b)\Bigr{)}$. By design, $\hat{h}_{l}$ is a sparse linear combinations of (latent) dictionary features.
3.2 Training
XNNTab is trained in four steps: i) training of the MLP’s representation, ii) training of the SAE, and iii) finetuning of the decision layer, iv) combining linear components.
Step 1: MLP training. The MLP’s parameters $\theta_{\text{MLP}}$= ($\theta_{g}$, $W$) are trained using cross-entropy loss and L1 regularization. The SAE is not present in this step.
Step 2: SAE Training. To learn a dictionary of sparse features we follow the training procedure by Huben et al. (2023). The SAE is trained to reconstruct the hidden activations $h_{l}$ of training samples $x$. The SAE uses a reconstructing loss and a sparsity loss on its hidden activations $h_{SAE}$ to learn sparse dictionary features.
Step 3: Finetuning the decision layer. We freeze the parameters $\theta_{g}$ of the MLP and $\theta_{SAE}$ of the SAE and finetune the weights $W$ in the decision layer with the same loss as in step 1 to make predictions from the reconstructions learned by the SAE $\hat{h}_{l}$.
[…]
3.3 Learning Representation Semantics
The hidden representations of the SAE represent monosemantic dictionary features in a latent space, but their semantics are unknown. To assign human-understandable semantics to those features, we follow a similar idea as in (Huben et al., 2023) for tabular data instead of text (cf. Figure 1, B). For each dimension in the latent representation of the SAE $j\in\{1,...d_{hid}\}$, we identify the subset of training samples $T_{j}$ which highly active this feature, i.e., whose activations are above a threshold $t$. We then use a rule-based classifier to learn simple decision rules describing this subset.
4 Experiments
Datasets and Evaluation Metrics. We evaluate on two benchmark datasets for structured data, Adult (ADULT(Vanschoren et al., 2014) and Churn Modelling (CHURN) [^3] https://www.kaggle.com/datasets/shrutimechlearn/churn-modelling. For ADULT, we only used the numerical features resulting in 6 features describing approx. 50k instances. CHURN contains 14 features and 10k instances. In both cases, the task is binary classification. For evaluation, we use 5-fold-cross-validation. The training/validation/test proportion of the datasets for each split are 65%/15%/20%. We report accuracy and macro F1 on the test set, averaged across all folds.
Models and Parameters. We used neural architecture search for the MLP and SAE components of XNNTab. The base MLP on ADULT has 3 layers with (97, 30, 7) neurons, the SAE $3\times 7=21$ (R=3) neurons in the hidden layer. For CHURN, the base MLP has 2 layers with (63, 24) neurons, the SAE $2\times 24=48$ (R=2) neurons in the hidden layer. See Appendix A.2 for details on all hyperparameters.
Representation Semantics. To assign human-understandable semantics to the dictionary features, we employ Skope-rules [^4] https://github.com/scikit-learn-contrib/skope-rules. Accessed July 2025. For each dictionary feature $f_{j}$, we retrieve a set $T_{j}$ of instances for which the activation of neuron $j$ is above the threshold $t=0.9$. We assign $label=1$ to instances in $T_{j}$ and $label=0$ to the remaining instances. The Skope-rules classifier learns a set of decision rules describing instances of the positive class. For each $T_{j}$, we keep the rule with the highest coverage (recall).
Results. Table 1 shows that, on both datasets, XNNTab outperforms interpretable models. On ADULT, XNNTab outperforms MLP and Random Forest and is comparable to XGBoost, on CHURN, XNNTab outperforms all baseline blackbox models including XGBoost. These results show that our model introduces interpretability to DNNs with very little compromise on performance. Rules extracted for the dictionary features on ADULT are shown in Table 2, for CHURN in Appendix B.2. The learned rules provide a global explanation of the model’s behavior (Figure 2 left), since the predictions are simply a linear combination of the learned features (rules). Figure 2, right shows that on average 8 out of the 21 features that we learn for ADULT are active. That means, one local explanation requires us to inspect 8 rules to understand the model’s prediction. However, the rules are simple to understand (two conditions long on average). The number of active features on average increase for CHURN but not the complexity of the rules (average rule length remains 2).
Table 1: Performance of blackbox (✕) and fully interpretable (✓) models on two benchmark datasets. Best values are marked bold, best values for interpretable models bold italic.
ADULT CHURN
F1-Macro Acc F1-Macro Acc
Random Forest 0.717 $\pm$ 0.003 0.806 $\pm$ 0.002 0.738 $\pm$ 0.002 0.857 $\pm$ 0.001
XGBoost 0.750 $\pm$ 0.006 0.843 $\pm$ 0.003 0.730 $\pm$ 0.000 0.854 $\pm$ 0.000
MLP 0.730 $\pm$ 0.004 0.824 $\pm$ 0.002 0.745 $\pm$ 0.001 0.857 $\pm$ 0.001
Logistic Regression 0.691 $\pm$ 0.000 0.815 $\pm$ 0.002 0.601 $\pm$ 0.006 0.808 $\pm$ 0.000
Decision Tree 0.709 $\pm$ 0.000 0.804 $\pm$ 0.000 0.718 $\pm$ 0.011 0.836 $\pm$ 0.009
XNNTab (ours) 0.733 $\pm$ 0.006 0.823 $\pm$ 0.002 0.759 $\pm$ 0.001 0.861 $\pm$ 0.005
Table 2: Dictionary features for the ADULT dataset. $|T_{j}|$ - size of the training subset that strongly activate feature $j$. Coverage of the rule reported as number of samples and percentage of samples. Table sorted by $|T_{j}|$.
j $|T_{j}|$ Description Coverage
5 6755 age $\leq 26$ and edu_num $\leq 12$ 4940 / 0.73
11 6217 age $\leq 26$ and cg $\leq 2969$ and edu_num $\leq 11$ 4722 / 0.76
16 3863 age $\leq 22$ and cg $\leq 3389$ and hpw $\leq 43$ 2604 / 0.67
15 3328 age $\leq$ and hpw $\leq$ 30.5 2082 / 0.62
9 2872 age $\leq 24$ and hpw $\leq 31$ 1919 / 0.66
21 1123 cg > 8296 and hpw > 22 751 / 0.68
6 789 age > 35 and cg > 9474 700 / 0.79
20 660 age $\leq 21$ and edu_num $\leq 9$ and cg $\leq 60025$ and hpw $\leq 23$ 406 / 0.61
1 210 cg > 22587.5 210 / 1.00
8 210 cg > 22588 and cl $\leq 78$ 210 / 1.00
13 203 age > 19 and cg > 25180 202 / 0.99
14 155 cg > 70655 and cl $\leq 78$ 155 / 1.00
ADULT CHURN
Rule Length
     min 1 1
     avg 2.2 1.8
     max 4 2
Active Features
     min 3 16
     avg 8 21.5
     max 11 28
Figure 2: Left: Decision weight matrix $W^{\prime}$ for ADULT on interpretable features. Right: Statistics on complexity of features on ADULT and CHURN.
Appendix A Implementation Details
A.1 Datasets
Table 3: Dataset overview
# Instances # Features Task
Adult 48 842 6 Binary classification
Churn Modeling 10 000 14 Binary classification
Data Details. ADULT contains demographic information such as age, education, and capital gain. The target variable is wether a person’s income is >50k or $\leq 50$. CHURN contains details of a bank’s customers, the target variable is whether the customer left the bank or continues being a customer.
Data preprocessing. We normalize the input features of ADULT. For CHURN, we drop the feature columns “RowNumber”, “CustomerId”, and “Surname”, encode the categorical columns and normalize the numerical ones to $[0,1]$ by min-max normalization.
A.2 Parameter settings
MLP. We use the Optuna library  [^5] https://optuna.org for neural architecture search and tune on the number of layers in the range of [1, 4], the dimension of the hidden layers (latent_dim) in the range of [1, 200], the learning rate ($lr$) in the range of [1e21e-2, 5e35e-3], and dropout rate ($Dropout$) in the range of [0.0, 0.5]. The parameter settings for ADULT are latent_dim = [97, 30, 7], Dropout = [0.3, 0.35, 0.23], lr=4e3lr=4e-3, and $\lambda=1e-4$ for L1 regularization. The parameter settings for CHURN are latent_dim = [63, 24], Dropout = [0.4, 0.3], lr=1e2lr=1e-2, and $\lambda=4e-4$ for L1 regularization.
XNNTab Follows the same setting details as MLP for the neural network component, the parameter settings for the sparse autoencoder are the $l1$ coefficient $\alpha=1e-3$ for ADULT and CHURN.
For the following baseline models we use the default parameter settings from scikit-learn (v1.7.1) [^6] https://scikit-learn.org/stable/
Random Forest. We tuned on number of estimators (n_estimators)(n\_estimators) from the list [100, 200].
XGBoost. We tuned on number of boosting iterations (num_boost_round)(num\_boost\_round) in the range of [100, 150, 200] and set the early stopping early_stopping_roundsearly\_stopping\_rounds to 12.
Decision Trees. We tuned on maximum depth (max_depth)(max\_depth) in the range of[5, 10, 15, 20].
Logistic Regression. We tuned on maximum of iterations (max_iter)(max\_iter) in the range of [100, 200].
For our rule-based classifier:
Skope-rules. we set precision_min=1precision\_min=1, recall_min=0.2recall\_min=0.2, and max_depth=4max\_depth=4, and the default settings for the rest of the parameters.
Appendix B Qualitative Results
B.1 ADULT
Table 4: Local explanation of an instance with positive prediction (> 50k) from ADULT. $c_{j}$ are the nonzero activations of dictionary features, the contribution of all activated features to the two output classes is $c^{\top}W^{\prime}$, where $W^{\prime}$ is the combined weight matrix.
Dataset features age fnlwgt edu_num cg cl hpw
Raw instance 48 175622 15 99999 0 60
Dictionary feature
$j$ rule $c_{j}$ $W^{\prime}_{j,0}$ $c_{j}\cdot W^{\prime}_{j,0}$ $W^{\prime}_{j,1}$ $c_{j}\cdot W^{\prime}_{j,1}$
1 cg > 22587.5 4.8735 -0.6028 -2.939 0.4310 2.101
6 age > 35 $\wedge$ cg > 9474 12.9912 -0.1803 -2.342 0.2962 3.848
8 cg > 22588 $\wedge$ cl $\leq 78$ 4.5567 0.1452 0.662 0.3139 1.431
9 age $\leq 24$ $\wedge$ hpw $\leq 31$ 5.1609 0.2681 1.384 -0.0763 -0.394
13 age > 19 $\wedge$ cg > 25180 3.9865 -0.6343 -2.528 0.4284 1.709
14 cg > 70655 $\wedge$ cl $\leq 78$ 1.8393 -0.1363 -0.251 -0.2378 -0.437
16 age $\leq 22$ $\wedge$ cg $\leq 3389$ $\wedge$ hpw $\leq 43$ 5.6880 0.1648 0.937 0.0393 0.224
20 age $\leq 21$ $\wedge$ edu_num $\leq 9$ $\wedge$ cg $\leq 60025$ $\wedge$ hpw $\leq 23$ 2.0693 0.1102 0.228 0.0218 0.045
21 cg > 8296 $\wedge$ hpw > 22 15.8202 -0.4210 -6.660 0.6348 10.039
Output Logits -11.509 18.567
Class $\mathbf{\leq 50K}$ $\mathbf{\geq 50K}$
Table 5: Local explanation of one instance with negative prediction ($\leq$ 50k) from ADULT. $c_{j}$ are the nonzero activations of dictionary features, the contribution of all activated features to the two output classes is $c^{\top}W^{\prime}$, where $W^{\prime}$ is the combined weight matrix. Dictionary features 7, 10 and 19 never exceeded the activation threshold $t=0.9$ and hence have no rules associated.
Dataset features Age fnlwgt edu_num cg cl hpw
Raw instance 18 225859 10 2907 0 30
Dictionary feature
$j$ rule $c_{j}$ $W^{\prime}_{j,0}$ $c_{j}\cdot W^{\prime}_{j,0}$ $W^{\prime}_{j,1}$ $c_{j}\cdot W^{\prime}_{j,1}$
5 age $\leq 26$ $\wedge$ edu_num $\leq 12$ 1.9102 0.5432 $\mathbf{1.0376}$ -0.6966 $\mathbf{-1.3306}$
6 age > 35 $\wedge$ cg > 9474 0.4034 -0.1803 $\mathbf{-0.0727}$ 0.2962 $\mathbf{0.1195}$
7 0.2820 0.1042 $\mathbf{0.0294}$ -0.5845 $\mathbf{-0.1648}$
9 age $\leq 24$ $\wedge$ hpw $\leq 31$ 0.9971 0.2681 $\mathbf{0.2673}$ -0.0763 $\mathbf{-0.0761}$
10 0.3919 0.3814 $\mathbf{0.1495}$ -0.3506 $\mathbf{-0.1374}$
11 age $\leq 26$ $\wedge$ cg $\leq 2969$ $\wedge$ edu_num $\leq 11$ 1.7615 0.3644 $\mathbf{0.6419}$ -0.5136 $\mathbf{-0.9047}$
14 cg > 70655 $\wedge$ cl $\leq 78$ 0.1123 -0.1363 $\mathbf{-0.0153}$ -0.2378 $\mathbf{-0.0267}$
15 age $\leq$ $\wedge$ hpw $\leq$ 30.5 1.1112 0.0624 $\mathbf{0.0693}$ -0.5594 $\mathbf{-0.6216}$
16 age $\leq 22$ $\wedge$ cg $\leq 3389$ $\wedge$ hpw $\leq 43$ 1.1850 0.1648 $\mathbf{0.1953}$ 0.0393 $\mathbf{0.0466}$
19 0.3856 0.3426 $\mathbf{0.1321}$ -0.0268 $\mathbf{-0.0103}$
20 age $\leq 21$ $\wedge$ edu_num $\leq 9$ $\wedge$ cg $\leq 60025$ $\wedge$ hpw $\leq 23$ 0.5693 0.1102 $\mathbf{0.0627}$ 0.0218 $\mathbf{0.0124}$
Output Logits 2.4662 -3.0938
Class $\mathbf{\leq 50K}$ $\mathbf{\geq 50K}$
B.2 CHURN
Table 6: Dictionary features for the CHURN dataset. $|T_{j}|$ - size of the training subset that strongly activate feature $j$. Coverage of the rule reported as number of samples and percentage of samples. Table sorted by $|T_{j}|$.
$j$ $|T_{j}|$ Description Coverage
6 4955 Age < 45 and NumOfProducts < 3 4560 / 0.91
47 4631 Age $\leq$ 55.5 and NumOfProducts $\geq$ 2 2898 / 0.62
2 4628 CreditScore > 379.0 and Age < 40 3561 / 0.76
3 4457 CreditScore > 409.0 and NumOfProducts $\geq$ 2 2939 /0.66
30 4419 Age < 45 and Balance > 76177.5 2594 /0.57
43 3966 Age $\leq$ 42 2238 / 0.62
9 2999 Age < 46.5 and NumOfProducts $\geq$ 2 2533 / 0.84
12 2574 Age < 35 and NumOfProducts $\geq$ 2 1193 / 0.46
18 1805 Geography_Spain == False and Age $\geq$ 45 1081 / 0.59
36 425 Age > 59.5 and IsActiveMember == 1 255 / 0.60
7 424 NumOfProducts $\geq$ 3 196 / 0.46
24 363 CreditScore > 361.5 and NumOfProducts $\geq$ 3 211 / 0.57
1 350 CreditScore > 379.0 and NumOfProducts $\geq$ 3 205 / 0.59
38 303 Age > 64.0 and IsActiveMember == 1 172 / 0.57
28 270 Balance $\leq$ 244642.5 and NumOfProducts $\geq$ 3 215 /0.79
40 226 NumOfProducts $\geq$ 3 196 / 0.87
11 218 Age > 62.5 and IsActiveMember == True 164 /0.74
35 150 Age $\geq$ 38 and Balance $\geq$ 75720 98 /0.65
18 146 Geography_Germany == True and IsActiveMember == False 54 /0.29
4 127 Age > 61.5 and NumOfProducts < 2 and IsActiveMember == True 101 / 0.79
16 112 Age > 66.0 and NumOfProducts < 2 and IsActiveMember == True 60 / 0.55
5 53 Age > 77 12 / 0.23

### v2

@@
 This allows us to represent predictions as linear combinations of semantically meaningful components.
-Empirical evaluations demonstrate that XNNTab attains performance on par with or exceeding that of state-of-the-art, black-box neural models and classical machine learning approaches while being fully interpretable.
+Empirical evaluations demonstrate that XNNTab attains performance comparable to that of state-of-the-art, black-box neural models and classical machine learning approaches while being fully interpretable.
 1 Introduction
@@
 To improve the performance of DNNs on tabular data, several deep learning methods have been developed with specialized architectures that take into account the unique traits of tabular data (mixture of categorical, ordinal and continuous features) in order to be on par (Huang et al., 2020) or even outperform GBDTs (Chen et al., 2024).
-While these methods addressed the limitations of DNNs in terms of performance, DNNs still remain blackboxes, whose decisions are hard to communicate to end users (Schwalbe and Finzel, 2023).
+While these methods address the limitations of DNNs in terms of performance, DNNs still remain blackboxes, whose decisions are hard to communicate to end users (Schwalbe and Finzel, 2023).
 In this paper, we address the interpretability limitation of existing DNNs and present XNNTab, a deep neural network with a sparse autoencoder (SAE) component.
@@
 Our method is illustrated in Figure 1 and described in Section 3.
-We show that XNNTab outperforms both interpretable and classical blackbox models and has similar performance to GBDTs.
-Additionally, XNNTab is fully interpretable: the final prediction is a simple linear combination of the interpretable dictionary features, which have easily accessible human-understandable semantics.
+We show that XNNTab outperforms interpretable models and has similar performance to classical blackbox models.
+Additionally, XNNTab is fully interpretable: the final prediction is a simple linear combination of the dictionary features, which have easily accessible human-understandable semantics.
 2 Related Work
 Gradient boosted decision trees (GBDTs) (Chen and Guestrin, 2016; Khan et al., 2022) have shown state-of-the art performance on tabular data (Grinsztajn et al., 2022).
-To close this performance gap to GBDTs, neural architectures specifically designed for tabular data have been proposed, such as attention-based architectures (Gorishniy et al., 2021; Somepalli et al., 2022; Yan et al., 2023), retrieval-augmented models (Somepalli et al., 2022; Gorishniy et al., 2023).
+To close this performance gap to GBDTs, neural architectures specifically designed for tabular data have been proposed, such as attention-based architectures (Gorishniy et al., 2021; Somepalli et al., 2022; Yan et al., 2023) or retrieval-augmented models (Somepalli et al., 2022; Gorishniy et al., 2023).
 Additional techniques include tailored regularizations (Jeffares et al., 2023), deep ensembles (Gorishniy et al., 2024), attentive feature selection strategies (Arik and Pfister, 2021), categorical feature embedding (Huang et al., 2020) and reconstruction of binned feature indices (Lee et al., 2024).
@@
 InterpreTabNet (Si et al., 2024), a variant of TabNet (Arik and Pfister, 2021) adds post-hoc explanations by learning sparse feature attribution masks and using LLMs to interpret the learned features from the masks.
-Similar to ProtoGate and different to InterpetTabNet, XNNTab is intrinsically interpretable, i.e., it uses interpretable features directly for prediction.
+Similar to ProtoGate and different from InterpetTabNet, XNNTab is intrinsically interpretable, i.e., it uses interpretable features directly for prediction.
 While ProtoGate relies on class prototypes, XNNTab learns relevant feature combinations that describe parts of instances, similar to part-prototype models (Elhadri et al., 2025).
@@
 3.1 Architecture
-The base blackbox model is an MLP and with multiple hidden layers and a softmax output.
+The base blackbox model is an MLP with multiple hidden layers and a softmax output.
 We denote the penultimate layer as $l$, and the representations of this layer as $h_{l}$.
@@
 $g(\cdot)$ is characterized by the parameters $\theta_{g}$, and $h(\cdot)$ is characterized by the weight matrix $W$.
-For the SAE, we use the architecture introduced by Huben et al.
+We adopt the SAE architecture introduced by Huben et al.
 (2023).
-The SAE consists of an input layer, a hidden layer and an output layer.
-The size of the hidden layer is $d_{hid}=R\cdot d_{in}$ with $d_{in}$ being the dimension of the MLP’s layer $l$, and $R$ a hyperparameter controlling the size of the feature dictionary.
+The SAE consists of an input layer, a hidden layer, and an output layer.
+The weights of the encoder and decoder are shared, and given by the matrices $M$ and $M^{T}$, respectively.
+The encoder has an additional bias vector $b$, the decoder has not (Huben et al., 2023).
+We set the size of the hidden layer as $d_{hid}=R\cdot d_{in}$ with $d_{in}$ being the dimension of the MLP’s layer $l$.
+$R$ is a hyperparameter that controls the size of the feature dictionary.
 The input to the SAE are the hidden representations of the blackbox MLP $h_{l}$.
-The weights of the encoder and decoder are tied, and given by the matrix $M$ ($M^{T}$ for the decoder)  [^1] Additionally, the encoder has a bias vector $b$..
 The output of the SAE is given by $\hat{h}_{l}=M^{T}\Bigl(ReLU(Mh_{l}+b)\Bigr)$.
-By design, $\hat{h}_{l}$ is a sparse linear combinations of (latent) dictionary features.
+By design, $\hat{h}_{l}$ is a sparse linear combination of (latent) dictionary features (Huben et al., 2023).
 3.2 Training
@@
 Step 2: SAE Training.
-To learn a dictionary of sparse features we follow the training procedure by Huben et al.
+To learn a dictionary of monosemantic features we follow the training procedure by Huben et al.
 (2023).
@@
 We evaluate on two benchmark datasets for structured data, Adult (ADULT(Vanschoren et al., 2014) and Churn Modelling (CHURN) [^2] https://www.kaggle.com/datasets/shrutimechlearn/churn-modelling.
-For ADULT, we only used the numerical features resulting in 6 features describing approx.
+ADULT contains 15 features describing approx.
 50k instances.
@@
 Models and Parameters.
-We used neural architecture search for the MLP and SAE components of XNNTab.
-The base MLP on ADULT has 3 layers with (97, 30, 7) neurons, the SAE $3\times 7=21$ (R=3) neurons in the hidden layer.
-For CHURN, the base MLP has 2 layers with (63, 24) neurons, the SAE $2\times 24=48$ (R=2) neurons in the hidden layer.
+We manually define the neural architecture for the MLP and SAE components of XNNTab.
+The base MLP has 3 layers with (100, 64, 32) neurons, the SAE $2\times 32=64$ (R=2) neurons in the hidden layer.
 See Appendix A.2 for details on all hyperparameters.
@@
 Accessed July 2025.
-For each dictionary feature $f_{j}$, we retrieve a set $T_{j}$ of instances for which the activation of neuron $j$ is above the threshold $t=0.9$.
+For each dictionary feature $f_{j}$, we retrieve a set $T_{j}$ of instances for which the activation of neuron $j$ is above the threshold $t=0.75$.
 We assign $label=1$ to instances in $T_{j}$ and $label=0$ to the remaining instances.
 The Skope-rules classifier learns a set of decision rules describing instances of the positive class.
+We set the precision hyperparameter of Skope-rules to 1 and keep the rest of the hyperparameter to the default settings.
 For each $T_{j}$, we keep the rule with the highest coverage (recall).
@@
 Table 1 shows that, on both datasets, XNNTab outperforms interpretable models.
-On ADULT, XNNTab outperforms MLP and Random Forest and is comparable to XGBoost, on CHURN, XNNTab outperforms all baseline blackbox models including XGBoost.
+On ADULT, XNNTab is comparable to MLP and slightly below the performance of XGBoost and Rando Forest.
+On CHURN, XNNTab is on par with all baseline blackbox models including XGBoost.
 These results show that our model introduces interpretability to DNNs with very little compromise on performance.
-Rules extracted for the dictionary features on ADULT are shown in Table 2, for CHURN in Appendix B.2.
+A selection of rules extracted for the dictionary features on ADULT are shown in Table 2 (full list in Appendix B.1, for CHURN in Appendix B.2).
 The learned rules provide a global explanation of the model’s behavior (Figure 2 left), since the predictions are simply a linear combination of the learned features (rules).
-Figure 2, right shows that on average 8 out of the 21 features that we learn for ADULT are active.
-That means, one local explanation requires us to inspect 8 rules to understand the model’s prediction.
-However, the rules are simple to understand (two conditions long on average).
-The number of active features on average increase for CHURN but not the complexity of the rules (average rule length remains 2).
+Each weight in this linear combination directly indicates whether the presence of the corresponding features increases or lowers the probability of a class, and to which extent.
+Figure 2 (right) shows that on average 30 out of the 64 features that we learn for ADULT are active above the chosen threshold.
+That means, one local explanation requires us to inspect 30 rules to understand the model’s prediction.
+The number of active features on average slightly increases for CHURN, but not the complexity of the rules (average rule length remains 2).
+While the number of active rules is high, several rules are redundant and could be pruned.
+A pruning mechanism could improve the interpretabiliy of the explanations, and while not in the scope of this paper, it is part of our future work.
 Table 1: Performance of blackbox (✕) and fully interpretable (✓) models on two benchmark datasets. Best values are marked bold, best values for interpretable models bold italic.
@@
 F1-Macro Acc F1-Macro Acc
-Random Forest 0.717 $\pm$ 0.003 0.806 $\pm$ 0.002 0.738 $\pm$ 0.002 0.857 $\pm$ 0.001
-XGBoost 0.750 $\pm$ 0.006 0.843 $\pm$ 0.003 0.730 $\pm$ 0.000 0.854 $\pm$ 0.000
-MLP 0.730 $\pm$ 0.004 0.824 $\pm$ 0.002 0.745 $\pm$ 0.001 0.857 $\pm$ 0.001
-Logistic Regression 0.691 $\pm$ 0.000 0.815 $\pm$ 0.002 0.601 $\pm$ 0.006 0.808 $\pm$ 0.000
-Decision Tree 0.709 $\pm$ 0.000 0.804 $\pm$ 0.000 0.718 $\pm$ 0.011 0.836 $\pm$ 0.009
-XNNTab (ours) 0.733 $\pm$ 0.006 0.823 $\pm$ 0.002 0.759 $\pm$ 0.001 0.861 $\pm$ 0.005
-Table 2: Dictionary features for the ADULT dataset.
+Random Forest 0.799 $\pm$ 0.003 0.861 $\pm$ 0.002 0.747 $\pm$ 0.013 0.862 $\pm$ 0.007
+XGBoost 0.815 $\pm$ 0.002 0.869 $\pm$ 0.001 0.750 $\pm$ 0.012 0.861 $\pm$ 0.007
+MLP 0.796 $\pm$ 0.002 0.850 $\pm$ 0.002 0.751 $\pm$ 0.106 0.860 $\pm$ 0.006
+Logistic Regression 0.782 $\pm$ 0.002 0.815 $\pm$ 0.002 0.601 $\pm$ 0.006 0.808 $\pm$ 0.000
+Decision Tree 0.787 $\pm$ 0.001 0.851 $\pm$ 0.002 0.735 $\pm$ 0.008 0.849 $\pm$ 0.006
+XNNTab (ours) 0.795 $\pm$ 0.002 0.850 $\pm$ 0.002 0.752 $\pm$ 0.008 0.861 $\pm$ 0.002
+Table 2: A selection of dictionary features for the ADULT dataset.
 $|T_{j}|$ - size of the training subset that strongly activate feature $j$.
@@
 j |T_j| Description Coverage
-5 6755 age $\leq 26$ and edu_num $\leq 12$ 4940 / 0.73
-11 6217 age $\leq 26$ and cg $\leq 2969$ and edu_num $\leq 11$ 4722 / 0.76
-16 3863 age $\leq 22$ and cg $\leq 3389$ and hpw $\leq 43$ 2604 / 0.67
-15 3328 age $\leq$ and hpw $\leq$ 30.5 2082 / 0.62
-9 2872 age $\leq 24$ and hpw $\leq 31$ 1919 / 0.66
-21 1123 cg > 8296 and hpw > 22 751 / 0.68
-6 789 age > 35 and cg > 9474 700 / 0.79
-20 660 age $\leq 21$ and edu_num $\leq 9$ and cg $\leq 60025$ and hpw $\leq 23$ 406 / 0.61
-1 210 cg > 22587.5 210 / 1.00
-8 210 cg > 22588 and cl $\leq 78$ 210 / 1.00
-13 203 age > 19 and cg > 25180 202 / 0.99
-14 155 cg > 70655 and cl $\leq 78$ 155 / 1.00
+40 15294 marital_status_Married is False and educational_num < 13 and capital_gain <= 8028.0 11625/0.76
+9 11968 marital_status_Married is False and age <= 37.5 and educational_num < 12 7271/0.61
+8 11634 marital_status_Married is False and age <= 34.5 and educational_num < 12 6517/0.56
+54 10889 marital_status_Married is False and age <= 31.5 and educational_num < 13 5953/0.55
+44 7457 marital_status_Married is False and relationship_Not_in_family is False and age <= 25.5 3330/0.45
+45 3975 marital_status_Married is False and age <= 22.5 and hours_per_week <= 32.5 1516/0.38
+5 1758 age <= 20.5 and hours_per_week <= 24.5 754/0.43
+34 1658 occupation_Other_service is False and capital_gain > 14682.0 547/0.33
+26 1358 marital_status_Widowedand is False and capital_gain > 14682.0 546/0.40
+50 982 age <= 72.0 and capital_gain > 15022.0 532/0.55
+63 978 gender is Female and age <= 18.5 and hours_per_week <= 24.5 236/0.24
+7 290 occupation_Farming_fishing is False and capital_gain > 19266.0 and hours_per_week > 27.5 209/0.73
+18 226 relationship_Other_relative is False and capital_gain > 26532.0 185/0.82
+22 225 occupation_Farming_fishing is False and capital_gain > 26532.0 and hours_per_week > 18.0 183/0.82
+3 212 capital_gain > 34569.0 153/0.73
 ADULT CHURN
@@
      min 1 1
-     avg 2.2 1.8
-     max 4 2
+     avg 2.18 1.8
+     max 3 2
 Active Features
-     min 3 16
-     avg 8 21.5
-     max 11 28
+     min 23 19
+     avg 30 33
+     max 37 40
 Figure 2: Left: Decision weight matrix $W^{\prime}$ for ADULT on interpretable features. Right: Statistics on complexity of features on ADULT and CHURN.
+Impact of activation threshold $t$.
+Figure 3 shows the impact of $t$ on the number of rules that can be extracted per active feature and on the number of instances covered by each rule (recall).
+The average recall increases with an increasing threshold, as the sets of instances by which features get activated become more focused.
+This makes it easier for Skope-rules to find a rule that explains most of the instances in the set.
+However, a focused set limits the data instances a rule could apply to.
+For some of these sets, Skope-rules fails to extract meaningful rules, i.e., the sets of instances that are activated by extracted rules become empty.
+In consequence, the amount of active features for which rules can be extracted decrease with increasing threshold $t$.
+Choosing the right threshold requires balancing the trade-off between the average rule recall and the fraction of extracted rules per active features.
+Figure 3: Fraction of features and instances covered by rules to explain ADULT at different feature activation thresholds $t$.
+Orange bars show the average fraction of rules extracted per active feature with the number of active features on top, blue bars show the average recall of the rules in terms of instances.
+5 Conclusion
+We introduced XNNTab, an interpretable model for tabular data.
+Our results show, that XNNTab outperforms interpretable models and performs comparably to blackbox architectures.
+In future work, we plan to investigate pruning of redundant rules and automated methods to identify activation thresholds for an optimal balance between rule recall and fraction of extracted rules.
+Funding Declaration
+This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under project ID 536124560.
 Appendix A Implementation Details
@@
 # Instances # Features Task
-Adult 48 842 6 Binary classification
+Adult 48 842 15 Binary classification
 Churn Modeling 10 000 14 Binary classification
@@
 A.2 Parameter settings
+For all the models, we use the Optuna library  [^4] https://optuna.org for hyperparameter tuning, the tuning is done on the validation set of each split (5 splits in total).
+For each experiment, the number of iterations is set to 100.
+Random seed = 0
 MLP.
-We use the Optuna library  [^5] https://optuna.org for neural architecture search and tune on the number of layers in the range of [1, 4], the dimension of the hidden layers (latent_dim) in the range of [1, 200], the learning rate ($lr$) in the range of [1e21e-2, 5e35e-3], and dropout rate ($Dropout$) in the range of [0.0, 0.5].
-The parameter settings for ADULT are latent_dim = [97, 30, 7], Dropout = [0.3, 0.35, 0.23], lr=4e3lr=4e-3, and $\lambda=1e-4$ for L1 regularization.
-The parameter settings for CHURN are latent_dim = [63, 24], Dropout = [0.4, 0.3], lr=1e2lr=1e-2, and $\lambda=4e-4$ for L1 regularization.
+We use a fixed architecture for both datasets [100, 64, 32].
+We tune on the learning rate ($lr$) in the range of [1e21e-2, 5e35e-3], dropout rate ($Dropout$) in the range of [0.0, 0.5] and Lambda ($\lambda$) in the range of [1e71e-7, 1e21e-2] for L1 regularization.
 XNNTab Follows the same setting details as MLP for the neural network component, the parameter settings for the sparse autoencoder are the $l1$ coefficient $\alpha=1e-3$ for ADULT and CHURN.
-For the following baseline models we use the default parameter settings from scikit-learn (v1.7.1) [^6] https://scikit-learn.org/stable/
-Random Forest. We tuned on number of estimators (n_estimators)(n\_estimators) from the list [100, 200].
-XGBoost. We tuned on number of boosting iterations (num_boost_round)(num\_boost\_round) in the range of [100, 150, 200] and set the early stopping early_stopping_roundsearly\_stopping\_rounds to 12.
-Decision Trees. We tuned on maximum depth (max_depth)(max\_depth) in the range of[5, 10, 15, 20].
+Random Forest.
+We tuned on number of estimators (n_estimators)(n\_estimators) from the list [100, 200, 300], maximum depth (max_depth)(max\_depth) in the range of [5, 10, 15, 20], the minimum samples split (min_samples_split)(min\_samples\_split) in the range of [2, 10], and the minimum samples leaf (min_samples_leaf)(min\_samples\_leaf) in the range of [1, 10].
+XGBoost.
+We set the booster to "gbtree", and "early_stopping_rounds" to 12 and tuned on learning rate (learning_rate)(learning\_rate) in the range of [0.001, 0.4], maximum depth $(max/_{d}epth)$ in the range of [3, 10], subsample in the range of [0.5, 1], and lambda $(\lambda)$ in the range of [0.1, 10]
+Decision Trees.
+We tuned on maximum depth (max_depth)(max\_depth) in the range of [5, 10, 15, 20], the minimum samples split (min_samples_split)(min\_samples\_split) in the range of [2, 10], and the minimum samples leaf (min_samples_leaf)(min\_samples\_leaf) in the range of [1, 10].
 Logistic Regression. We tuned on maximum of iterations (max_iter)(max\_iter) in the range of [100, 200].
 For our rule-based classifier:
-Skope-rules. we set precision_min=1precision\_min=1, recall_min=0.2recall\_min=0.2, and max_depth=4max\_depth=4, and the default settings for the rest of the parameters.
+Skope-rules. we set precision_min=1precision\_min=1, recall_min=0.2recall\_min=0.2, and the default settings for the rest of the parameters.
 Appendix B Qualitative Results
 B.1 ADULT
-Table 4: Local explanation of an instance with positive prediction (> 50k) from ADULT.
-$c_{j}$ are the nonzero activations of dictionary features, the contribution of all activated features to the two output classes is $c^{\top}W^{\prime}$, where $W^{\prime}$ is the combined weight matrix.
-Dataset features age fnlwgt edu_num cg cl hpw
-Raw instance 48 175622 15 99999 0 60
-Dictionary feature
-$j$ rule $c_{j}$ $W^{\prime}_{j,0}$ $c_{j}\cdot W^{\prime}_{j,0}$ $W^{\prime}_{j,1}$ $c_{j}\cdot W^{\prime}_{j,1}$
-1 cg > 22587.5 4.8735 -0.6028 -2.939 0.4310 2.101
-6 age > 35 $\wedge$ cg > 9474 12.9912 -0.1803 -2.342 0.2962 3.848
-8 cg > 22588 $\wedge$ cl $\leq 78$ 4.5567 0.1452 0.662 0.3139 1.431
-9 age $\leq 24$ $\wedge$ hpw $\leq 31$ 5.1609 0.2681 1.384 -0.0763 -0.394
-13 age > 19 $\wedge$ cg > 25180 3.9865 -0.6343 -2.528 0.4284 1.709
-14 cg > 70655 $\wedge$ cl $\leq 78$ 1.8393 -0.1363 -0.251 -0.2378 -0.437
-16 age $\leq 22$ $\wedge$ cg $\leq 3389$ $\wedge$ hpw $\leq 43$ 5.6880 0.1648 0.937 0.0393 0.224
-20 age $\leq 21$ $\wedge$ edu_num $\leq 9$ $\wedge$ cg $\leq 60025$ $\wedge$ hpw $\leq 23$ 2.0693 0.1102 0.228 0.0218 0.045
-21 cg > 8296 $\wedge$ hpw > 22 15.8202 -0.4210 -6.660 0.6348 10.039
-Output Logits -11.509 18.567
-Class $\mathbf{\leq 50K}$ $\mathbf{\geq 50K}$
-Table 5: Local explanation of one instance with negative prediction ($\leq$ 50k) from ADULT.
-$c_{j}$ are the nonzero activations of dictionary features, the contribution of all activated features to the two output classes is $c^{\top}W^{\prime}$, where $W^{\prime}$ is the combined weight matrix.
-Dictionary features 7, 10 and 19 never exceeded the activation threshold $t=0.9$ and hence have no rules associated.
-Dataset features Age fnlwgt edu_num cg cl hpw
-Raw instance 18 225859 10 2907 0 30
-Dictionary feature
-$j$ rule $c_{j}$ $W^{\prime}_{j,0}$ $c_{j}\cdot W^{\prime}_{j,0}$ $W^{\prime}_{j,1}$ $c_{j}\cdot W^{\prime}_{j,1}$
-5 age $\leq 26$ $\wedge$ edu_num $\leq 12$ 1.9102 0.5432 $\mathbf{1.0376}$ -0.6966 $\mathbf{-1.3306}$
-6 age > 35 $\wedge$ cg > 9474 0.4034 -0.1803 $\mathbf{-0.0727}$ 0.2962 $\mathbf{0.1195}$
-7 0.2820 0.1042 $\mathbf{0.0294}$ -0.5845 $\mathbf{-0.1648}$
-9 age $\leq 24$ $\wedge$ hpw $\leq 31$ 0.9971 0.2681 $\mathbf{0.2673}$ -0.0763 $\mathbf{-0.0761}$
-10 0.3919 0.3814 $\mathbf{0.1495}$ -0.3506 $\mathbf{-0.1374}$
-11 age $\leq 26$ $\wedge$ cg $\leq 2969$ $\wedge$ edu_num $\leq 11$ 1.7615 0.3644 $\mathbf{0.6419}$ -0.5136 $\mathbf{-0.9047}$
-14 cg > 70655 $\wedge$ cl $\leq 78$ 0.1123 -0.1363 $\mathbf{-0.0153}$ -0.2378 $\mathbf{-0.0267}$
-15 age $\leq$ $\wedge$ hpw $\leq$ 30.5 1.1112 0.0624 $\mathbf{0.0693}$ -0.5594 $\mathbf{-0.6216}$
-16 age $\leq 22$ $\wedge$ cg $\leq 3389$ $\wedge$ hpw $\leq 43$ 1.1850 0.1648 $\mathbf{0.1953}$ 0.0393 $\mathbf{0.0466}$
-19 0.3856 0.3426 $\mathbf{0.1321}$ -0.0268 $\mathbf{-0.0103}$
-20 age $\leq 21$ $\wedge$ edu_num $\leq 9$ $\wedge$ cg $\leq 60025$ $\wedge$ hpw $\leq 23$ 0.5693 0.1102 $\mathbf{0.0627}$ 0.0218 $\mathbf{0.0124}$
-Output Logits 2.4662 -3.0938
-Class $\mathbf{\leq 50K}$ $\mathbf{\geq 50K}$
+Table 4: Dictionary features for the ADULT dataset.
+$|T_{j}|$ - size of the training subset that strongly activate feature $j$.
+Coverage of the rule reported as number of samples and percentage of samples.
+Table sorted by $|T_{j}|$.
+j |T_j| Description Coverage
+40 15294 marital_status_Married is False and educational_num < 13 and capital_gain <= 8028.0 11625/0.76
+9 11968 marital_status_Married is False and age <= 37.5 and educational_num < 12 7271/0.61
+8 11634 marital_status_Married is False and age <= 34.5 and educational_num < 12 6517/0.56
+54 10889 marital_status_Married is False and age <= 31.5 and educational_num < 13 5953/0.55
+37 10755 marital_status_Married is False and age <= 31.5 and educational_num < 12 5735/0.53
+49 9304 marital_status_Married is False and age <= 29.5 and educational_num < 11 4920/0.52
+11 9030 marital_status_Married is False and age <= 26.5 and educational_num < 13 4387/0.49
+31 8385 marital_status_Married is False and age <= 26.5 and educational_num < 11 4098/0.49
+38 8313 marital_status_Married is False and age <= 25.5 and educational_num < 13 4053/0.48
+44 7457 marital_status_Married is False and relationship_Not_in_family is False and age <= 25.5 3330/0.45
+45 3975 marital_status_Married is False and age <= 22.5 and hours_per_week <= 32.5 1516/0.38
+5 1758 age <= 20.5 and hours_per_week <= 24.5 754/0.43
+34 1658 occupation_Other_service is False and capital_gain > 14682.0 547/0.33
+26 1358 marital_status_Widowedand is False and capital_gain > 14682.0 546/0.40
+50 982 age <= 72.0 and capital_gain > 15022.0 532/0.55
+63 978 gender is Female and age <= 18.5 and hours_per_week <= 24.5 236/0.24
+12 521 occupation_Handlers_cleaners is False and educational_num > 10 and capital_gain > 15022.0 404/0.79
+61 480 capital_gain > 70654.5 152/0.32
+7 290 occupation_Farming_fishing is False and capital_gain > 19266.0 and hours_per_week > 27.5 209/0.73
+18 226 relationship_Other_relative is False and capital_gain > 26532.0 185/0.82
+22 225 occupation_Farming_fishing is False and capital_gain > 26532.0 and hours_per_week > 18.0 183/0.82
+3 212 capital_gain > 34569.0 153/0.73
+58 154 capital_gain > 37702.5 153/0.99
+42 153 capital_gain > 37702.5 153/1.00
+36 152 capital_gain > 70654.5 152/1.00
+4 152 capital_gain > 70654.5 152/1.00
+33 152 capital_gain > 67047.0 152/1.00
+53 152 capital_gain > 70654.5 152/1.00
+21 152 capital_gain > 70654.5 152/1.00
+60 152 capital_gain > 70654.5 152/1.00
+16 152 capital_gain > 70654.5 152/1.00
+13 152 capital_gain > 70654.5 152/1.00
+64 152 capital_gain > 70654.5 152/1.00
 B.2 CHURN